OpenStudy (anonymous):

fan and medal Solve for x: |x − 2| + 10 = 12 x = 0 and x = 4 x = −4 and x = 0 x = −20 and x = 4 No solution

OpenStudy (anonymous):

|5x+6|=41 |a|=b implies a=b and a=-b so 5x+6 = 41 and 5x+6 = -41 solve both of these equations for x x = 7, and x = -47/5 but what is the answer

OpenStudy (anonymous):

x = −4 and x = 0

OpenStudy (anonymous):

x = −4 and x = 0

OpenStudy (cggurumanjunath):

in 5x+6 there r 2 values for x one is 7 and other is -47/5

OpenStudy (jdoe0001):

solve both of these equations for x x = 7, and x = -47/5 <--------------- those ones but what is the answer for an absolute value expression, you'd always get 2 values for "x" or whatever variable

OpenStudy (cggurumanjunath):

similarly try urself for your first posted q . @sunshinee

OpenStudy (anonymous):

i didint want to solve anything , i just wanted the answer

OpenStudy (jdoe0001):

the idea behind exercises isn't to have folks transcribing material over is to actually solve them thus understanding them otherwise the exercises are meaningless

OpenStudy (anonymous):

algebra is meaningless i juss wanna graduate and try and get into the army